You have a stock of streptomycin (antibiotic) with a concentration of 200 mg/mL. You add water until the concentration of streptomycin is 8 mg/mL. What dilution have you performed? (Hint: Think of how many times less concentrated your final concentration is compared to your initial. Using C1V1=C2V2) This is all the information provided in the question.
Answer #1

C1V1 = C2V2

Now, initial concentration (C1) = 200 mg/mL

final concentration (C2) = 8 mg/mL

V1 are V2 the initial and final volumes respectively;

From C1V1 = C2V2

   or V1 / V2 = C2 / C1

or V1 / V2 = 8 / 200 = 1 / 25

So,   V1 : V2 = 1 : 25

So the final solution will be 25 times less concentrated than the initial solution.